OpenStudy (anonymous):

D/14-9=5 I NEED TO KNOW WHAT D EQUALS

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

myininaya (myininaya):

add 9 to both sides D/14-9+9=5+9 D/14+0=14 D/14=14 multiply 14 on both sides 14*D/14=14*14 D=14^2

OpenStudy (anonymous):

what does ^ mean

OpenStudy (anonymous):

whut about dhis question a/35+20=18

myininaya (myininaya):

subtract 20 on both sides a/35+20-20=18-20 a/35+0=-2 a/35=-2 multipliy 35 on both sides 35*a/35=-2*35 a=-2*35

OpenStudy (anonymous):

last one 3/4-9p=-3/5

myininaya (myininaya):

you need to get p by itself first step get -9p by itself-> ask yourself how can you do this?

OpenStudy (anonymous):

i dont kno how dhats y iym asking you

OpenStudy (anonymous):

so cute

OpenStudy (anonymous):

who

myininaya (myininaya):

what can you subtract on both sides that will give you -9p on the left side?